# LLM-based relevance assessment still can’t replace human relevance assessment

### v1

@@ Abstract
The use of large language models (LLMs) for relevance assessment in information retrieval has gained significant attention, with recent studies suggesting that LLM-based judgments provide comparable evaluations to human judgments. Notably, based on TREC 2024 data, Upadhyay et al. [8] make a bold claim that LLM-based relevance assessments, such as those generated by the UMBRELA system, can fully replace traditional human relevance assessments in TREC-style evaluations. This paper critically examines this claim, highlighting practical and theoretical limitations that undermine the validity of this conclusion. First, we question whether the evidence provided by Upadhyay et al. [8] really supports their claim, particularly if a test collection is used as a benchmark for future improvements. Second, through a submission deliberately intended to do so, we demonstrate the ease with which automatic evaluation metrics can be subverted, showing that systems designed to exploit these evaluations can achieve artificially high scores. Theoretical challenges – such as the inherent narcissism of LLMs, the risk of overfitting to LLM-based metrics, and the potential degradation of future LLM performance – must be addressed before LLM-based relevance assessments can be considered a viable replacement for human judgments.
1 Introduction
@@
Almost two years ago, Faggioli et al. [5] applied GPT-3.5 (text-davinci-003) to fully reassess runs submitted to the TREC 2021 Deep Learning track. They reported Kendall’s $\tau=0.86$ for human vs. LLM-based assessment on NDCG@10. A natural conclusion might be that LLMs could now replace humans for routine relevance assessment. Instead, Faggioli et al. [5] issue a warning. While recognizing the potential of LLMs to improve ranking and acknowledging their value as part of the relevance assessment process, they argue strongly against abandoning human assessment. They raise concerns about the potential for unknown biases that LLM-based assessments might introduce. They highlight the issue of circularity, where LLMs evaluate the outputs of other LLMs. Most importantly, their primary concern is that “LLMs are not people.” Since information retrieval systems are designed to serve human needs, their evaluation must ultimately reflect human judgment and preferences.
@@
Recently, Upadhyay et al. [8] analyze the retrieval task results from the TREC 2024 RAG track. This retrieval task (or “R task”) mirrors a traditional TREC ad hoc retrieval task. Participating systems were tasked with executing 301 queries over the MS MARCO Segment V2.1 collection, producing a ranked set of 100 passages for each query (a “run”). They compare four procedures for assessing runs: 1) fully automatic assessments using the UMBRELA LLM-based relevance assessment tool [9]; 2) fully manual assessments using established TREC evaluation protocols for human judgments; 3) a hybrid method where UMBRELA filtered the set of passages to be judged; and 4) a hybrid method where humans refined UMBRELA’s assessments. In this paper, we focus on the first two procedures: fully automatic and fully manual assessments. Based on their analysis, Upadhyay et al. [8] conclude:
@@
We disagree. Not only does their reported correlation fail to provide stronger evidence than that of Faggioli et al. [5], but additional evidence from the track directly contradicts their conclusion. This evidence includes runs submitted by one of the authors’ research teams (WaterlooClarke), which were explicitly designed to subvert LLM-based relevance judgments by employing LLM-generated judgments as a final-stage ranker.
Faggioli et al. [5] already demonstrated a strong empirical correlation between manual judgments and LLM judgments, both in terms of inter-annotator agreement and leaderboard correlation, and many other work also observed this empirical correlation. However, after a detailed consideration of competing views, Faggioli et al. [5] concluded that there are too many theoretical concerns before human judgments can be replaced. These concerns, which remain critical to the discussion, have neither been addressed nor refuted in the work of Upadhyay et al. [8, 9].
Conceptually, there is no fundamental difference between an LLM-based relevance assessment and an LLM-based re-ranking method. Human relevance judgments are privileged precisely because they are created by humans, and only humans can provide a gold standard for the evaluation of usefulness. While employing LLMs to train and implement rankers can lead to substantial performance gains, these improvements risk being illusory if they fail to reflect human judgments. The observation that LLM-based relevance judgments closely mimic the outcomes of human relevance judgments suggests that these LLM assessments may themselves represent a strong ranking method, rather than a valid evaluation metric.
We acknowledge the value of work by Upadhyay et al. [9]. In particular, their use of a new collection with fresh queries, which guarantees that LLMs were not trained on this collection. Moreover, their work confirms that the correlation seen by Faggioli et al. [5] is not merely an artifact of training on the test collection. However, the warnings issued by Faggioli et al. [5] remain both valid and increasingly urgent, especially given the growing prevalence of LLM-based relevance assessments in information retrieval tasks.
@@
Figure 1: Scatter plot extracted from Figure 3 of Upadhyay et al. [8] comparing manual assessment and automatic assessment. Each red dot represents the average performance of a run over all queries. The blue dots plot all run-query combination. The inset provides a closer view of the points in the green rectangle.
2 Differences among Top-Performing Systems
While Upadhyay et al. [8] report an overall Kendall’s $\tau=0.89$ for manual vs. automatic assessment, the correlation appears much weaker on the top-scoring runs, where agreement with manual judgments matters most. Although Kendall’s $\tau$ is a suitable metric for showing overall rank correlations on large-scale experiments, it is less informative for assessing progress at the top of the leaderboard. Demonstrating methodological advancements often involves reusing test collections with the goal of surpassing the best-performing systems. Therefore, identifying meaningful differences at the top of the leaderboard is critical for measuring significant progress. Based on the results reported by Upadhyay et al. [9], automatic judgments fail to demonstrate strong alignment with manual judgments for top-performing systems. This misalignment undermines the reliability of automatic assessments in distinguishing improvements among top performing systems.
@@ 2 Differences among Top-Performing Systems
Closer consideration of results from Upadhyay et al. [8] highlights specific discrepancies in top-performing system rankings. Figure 1 reproduces a scatter plot extracted from Figure 3 of that paper, showing the performance of submitted runs (red dots) under manual assessment (y-axis) vs. automatic assessment (x-axis).111 At the time of writing, Upadhyay et al. [8] do not provide a full public data release. They have generously provided limited access to their data for the purpose of confirming factual statements in this paper. Focusing on the top-performing systems in the inset of Figure 1, discrepancies between manual and automatic evaluations become evident. For example, the system ranked highest under automatic evaluation would only place fifth under manual evaluation. Conversely, the top system under manual evaluation ranks sixth under automatic evaluation. A particularly interesting is the case is the run circled in green. While it ranks fifth under automatic evaluation, it drops dramatically to 28th under manual evaluation.
@@
These inconsistencies underscore a fundamental limitation of LLM-based assessments, such as those used by UMBRELA, in reliably identifying the best-performing systems. As a result, caution is warranted when using these methods for evaluating and validating progress in retrieval tasks. While these differences may be the attributed to statistical noise — only 27 queries were manually judged, and no error bars are provided — the evidence presented is insufficient to justify the replacement of human judgments. This concern is particularly important when a collection is intended for re-use. To substantiate claims of a novel system’s superiority over state-of-the-art methods, we must ensure that any observed improvements, measured through fully automatic judgments, are both meaningful and aligned with manual assessments. At present, such confidence is lacking.
3 Subverting Automatic Evaluation
@@
When task relevance labels are generated entirely through a publicly known automatic process, such as UMBRELA, the evaluation metric becomes vulnerable to manipulation. For instance, a participant could pool the output of many rankers, apply the UMBRELA system to this pool, and submit the resulting relevance labels as a participating system for evaluation. This approach could theoretically achieve perfect scores across all metrics. Even if the specific LLM-based relevance assessment process includes undisclosed elements, such as the exact prompt or LLM used, participants could approximate the process to subvert automatic evaluations.
3.1 Empirical Demonstration of the Risk
@@
The run circled in green in Figure 1 exemplifies this approach. Submitted by Clarke’s research group (WaterlooClarke) as run uwc1, this submission was deliberately designed to subvert the automatic evaluation process. They pooled the top 20 documents from 15 preliminary runs, spanning neural and traditional rankers, with and without query expansion. This pool was then judged by GPT-4o with the prompt described by Faggioli et al. [5]. Top-graded passages were subsequently judged pairwise following the procedure of Clarke et al. [3] but using LLM-based assessments, rather than crowdsourced human judgments.
@@
The final ranking for uwc1 was determined using LLM-based preference judgments as the primary key, LLM-based relevance assessments as the secondary key, and the reciprocal rank fusion [4] of preliminary runs as the tertiary key. As intended, this deliberate attempt to manipulate the evaluation process led uwc1 to rank significantly higher under automatic assessment (5th) than under manual assessment (28th).
3.2 LLM-based Relevance Assessment as Re-Ranking
@@
LLM-based relevance assessment is effectively just another form of LLM-based re-ranking. To employ an LLM-based relevance assessment tool as a LLM-based re-ranker, one starts with an initial ranking and we prompt an LLM to assign a score — expressed as a relevance grade — to each of its top-$k$ passages. These passages are then re-ranked according to their relevance grades, preserving their original order when grades are equal.
@@
This concept is further illustrated by another run from WaterlooClarke, uwc2, which demonstrates that LLM-based relevance assessment operates as a pointwise re-ranking method. For uwc2, the group re-ranked the track’s baseline run using the prompt described by Arabzadeh and Clarke [2]. This re-ranking improves the baseline’s performance from 9th to 4th under manual assessments and from 7th to 3rd under automatic assessments.
4 Automatic Relevance Labels are not Gold Standards
@@
LLM-based relevance assessments cannot serve as a gold standard because they lack the grounding in human judgment necessary to evaluate the usefulness of retrieval systems. A true gold standard must originate from human assessments, as only humans can determine the relevance of information in a way that reflects real-world utility.
@@
Faggioli et al. [5] raised concerns about the potential unknown biases inherent in LLM-based assessments. However, one clear and concerning bias is that LLM-based relevance assessments tend to favor LLM-based ranking systems. This bias has been observed in other contexts as well [6, 10, 7], where LLMs demonstrate a form of “narcissism,” disproportionately favoring outputs generated by similar models. Furthermore, Alaofi et al. [1] show that LLMs can be deceived through well-crafted prompt attacks embedded in content, leading them to incorrectly judge irrelevant text as relevant. These vulnerabilities highlight not only the susceptibility of LLM-based assessments to manipulation but also their inability to objectively evaluate diverse ranking approaches. Such biases and flaws further undermine the reliability of LLM-based assessments as a substitute for human judgments in critical tasks.
5 When Automatic Judgments become Useless

### v2

@@ Abstract.
The use of large language models (LLMs) for relevance assessment in information retrieval has gained significant attention, with recent studies suggesting that LLM-based judgments provide comparable evaluations to human judgments. Notably, based on TREC 2024 data, Upadhyay et al. (2024a) make a bold claim that LLM-based relevance assessments, such as those generated by the Umbrela system, can fully replace traditional human relevance assessments in TREC-style evaluations. This paper critically examines this claim, highlighting practical and theoretical limitations that undermine the validity of this conclusion.
First, we question whether the evidence provided by Upadhyay et al. (2024a) genuinely supports their claim, particularly when the test collection is intended to serve as a benchmark for future research innovations. Second, we submit a system deliberately crafted to exploit automatic evaluation metrics, demonstrating that it can achieve artificially inflated scores without truly improving retrieval quality. Third, we simulate the consequences of circularity by analyzing Kendall’s tau correlations under the hypothetical scenario in which all systems adopt Umbrela as a final-stage re-ranker, illustrating how reliance on LLM-based assessments can distort system rankings. Theoretical challenges — including the inherent narcissism of LLMs, the risk of overfitting to LLM-based metrics, and the potential degradation of future LLM performance — that must be addressed before LLM-based relevance assessments can be considered a viable replacement for human judgments.
1. Introduction
@@
In early 2023, Faggioli et al. (2023) applied GPT-3.5 (text-davinci-003) to fully reassess runs submitted to the TREC 2021 Deep Learning track. They reported Kendall’s $\tau=0.86$ for human vs. LLM-based assessment on NDCG@10. A natural conclusion might be that LLMs could now replace humans for routine relevance assessment. Instead, Faggioli et al. (2023) issue a warning. While recognizing the potential of LLMs to improve ranking and acknowledging their value as part of the relevance assessment process, they argue strongly against abandoning human assessment. They raise concerns about the potential for unknown biases that LLM-based assessments might introduce. They highlight the issue of circularity, where LLMs evaluate the outputs of other LLMs. Most importantly, their primary concern is that “LLMs are not people.” Since information retrieval systems are designed to serve human needs, their evaluation must ultimately reflect human judgment and preferences.
@@
Recently, Upadhyay et al. (2024a) analyze the retrieval task results from the TREC 2024 RAG track. This retrieval task (or “R task”) mirrors a traditional TREC ad hoc retrieval task. Participating systems were tasked with executing 301 queries over the MS MARCO Segment V2.1 collection, producing a ranked set of 100 passages for each query (a “run”). They compare four procedures for assessing runs: 1) fully automatic assessments using the Umbrela LLM-based relevance assessment tool (Upadhyay et al., 2024b); 2) fully manual assessments using established TREC evaluation protocols for human judgments; 3) a hybrid method where Umbrela filtered the set of passages to be judged; and 4) a hybrid method where humans refined Umbrela’s assessments. In this paper, we focus on the first two procedures: fully automatic and fully manual assessments. Based on their analysis, Upadhyay et al. (2024a) conclude:
@@
We disagree. Not only does their reported correlation fail to provide stronger evidence than that of Faggioli et al. (2023), but additional evidence from the track directly contradicts their conclusion. This evidence includes runs submitted by team WaterlooClarke, which were explicitly designed to subvert LLM-based relevance judgments by employing LLM-generated judgments as a final-stage ranker.
Faggioli et al. (2023) already demonstrated a strong empirical correlation between manual judgments and LLM judgments, both in terms of inter-annotator agreement and leaderboard correlation, and many other work also observed this empirical correlation. However, after a detailed consideration of competing views, Faggioli et al. (2023) concluded that there are too many theoretical concerns before human judgments can be replaced. These concerns, which remain critical to the discussion, have neither been addressed nor refuted in the work of Upadhyay et al. (2024a, b).
Conceptually, there is no fundamental difference between an LLM-based relevance assessment and an LLM-based re-ranking method. Both predict an affinity score for a passage to be relevant for a given query. In contrast, human relevance judgments are privileged precisely because they are created by humans, and only humans can provide a gold standard for the evaluation of usefulness. While employing LLMs to train and implement rankers can lead to substantial performance gains, these improvements risk being illusory if they fail to reflect human judgments. The observation that LLM-based relevance judgments closely mimic the outcomes of human relevance judgments suggests that these LLM assessments may themselves represent a strong ranking method, rather than a valid evaluation metric.
We acknowledge the value of work by Upadhyay et al. (2024b). In particular, their use of a new collection with fresh queries, which guarantees that LLMs were not trained on this collection. Moreover, their work confirms that the correlation seen by Faggioli et al. (2023) is not merely an artifact of training on the test collection. However, the warnings issued by Faggioli et al. (2023) remain both valid and increasingly urgent, especially given the growing prevalence of LLM-based relevance assessments in information retrieval tasks.
@@
Figure 1. Scatter plot extracted from Figure 3 of Upadhyay et al. (2024a) comparing manual assessment and automatic assessment. Each red dot represents the average performance of a run over all queries. The blue dots plot all run-query combination. The inset provides a closer view of the points in the green rectangle.
2. Reproduction of Umbrela Results
Kendall’s $\tau$ is a suitable metric for showing overall rank correlations on large-scale experiments. Upadhyay et al. (2024a) report a high overall Kendall’s $\tau=0.89$ between manual relevance assessments and automatic Umbrela assessments. However, we note that some submitted systems perform substantially worse than others, making them easy to distinguish by an evaluation. The presence of such under-performing systems can artificially inflate Kendall’s $\tau$ scores.
To investigate this effect, we extend the original analysis by excluding the bottom-ranked 15 systems (out of 75) and recomputing Kendall’s $\tau$ over the remaining top 60 systems. As shown in Figure 2, this yields a slightly lower but still relatively high Kendall’s $\tau$ of 0.84. This corresponds to approximately 8% of system swaps, where the two evaluators disagree on which system performs better. Overall, these findings broadly confirm the results reported by Upadhyay et al. (2024a). Nevertheless, we observe some notable outliers. For example, one system achieves a high LLM-based evaluation score (0.81) but a substantially lower manual evaluation score (0.63), as shown in Figure 2. We analyze such discrepancies further in Sections 4.1 and 4.2.
Figure 2. Reproduction of Upadhyay et al. (2024a): On top 60 original TREC RAG 24 systems and data, the Umbrela LLM evaluator correlates highly with manual assessors. Only few submitted retrieval systems included approaches from LLM evaluators. Each system represents one dot. Red dots mark systems provided by team WaterlooClarke, which are known to contain LLM evaluations for re-ranking.
3. Differences among Top-Performing Systems
Demonstrating methodological advancements often involves reusing test collections with the goal of surpassing state-of-the-art systems. Identifying meaningful differences among top-performing runs is therefore critical for measuring significant progress. In contrast to the overall Kendall’s $\tau$ of 0.89 between manual and automatic assessments, the correlation weakens substantially among the highest-scoring systems. While Kendall’s $\tau$ among the top 60 systems [^1] Throughout the paper, when we refer to “top-performing systems”, we use the manual relevance judgments as a basis for this assessment. remains relatively high at 0.85 (corresponding to 8% system swaps), it drops to 0.51 (24% swaps) among the top 20 systems, and to 0.56 (21% swaps) among the top 15.
[…]
@@ 3. Differences among Top-Performing Systems
Closer consideration of results from Upadhyay et al. (2024a) further highlights specific discrepancies in top-performing system rankings. Figure 1 reproduces a scatter plot extracted from Figure 3 of that paper, showing the performance of submitted runs (red dots) under manual assessment (y-axis) vs. automatic assessment (x-axis).222 At the time of writing, Upadhyay et al. (2024a) do not provide a full public data release. They have generously provided limited access to their data for the purpose of confirming factual statements in this paper. Focusing on the top-performing systems (inset of Figure1), discrepancies become evident: for instance, the system ranked highest under automatic evaluation would only place fifth under manual evaluation. Conversely, the top system under manual evaluation ranks sixth under automatic evaluation. Particularly interesting is the case of the run circled in green. While it ranks fifth under automatic evaluation, it drops to 28th under manual evaluation.
@@
These inconsistencies underscore a fundamental limitation of LLM-based assessments, such as those used by Umbrela, in reliably identifying the best-performing systems. As a result, caution is warranted when using these methods for evaluating and validating progress in retrieval tasks. While some differences could be the attributed to statistical noise — only 27 queries were manually judged, and no error bars are provided — the available evidence remains insufficient to justify the replacement of human judgments. This concern is particularly important when a collection is intended for re-use: claims of a novel system’s superiority over existing methods must be supported by improvements that are both statistically meaningful and aligned with manual assessments. At present, such confidence is lacking.
Figure 3. Reranking with an LLM evaluator (Umbrela) improves performance under human relevance labels. This plot compares the original and reranked versions of all TREC RAG 24 systems based on manual assessment.
4. Subverting Automatic Evaluation
@@
When task relevance labels are generated entirely through a publicly known automatic process, such as Umbrela, the evaluation metric becomes vulnerable to manipulation. For instance, a participant could aggregate the outputs of many rankers, apply the Umbrela system to this pooled set, and submit the resulting relevance labels as a new system for evaluation. Such a strategy could, in principle, achieve perfect scores across all metrics. Even if the specific LLM-based relevance assessment process includes undisclosed elements, such as the exact prompt or LLM used, participants could approximate the process enough to subvert the automatic evaluations.
Figure 4. Demonstration of the effects of circularity when using Umbrela as both evluator and ranker using TREC RAG 24 data. Each submitted retrieval system is first re-ranked with Umbrela, then evaluated under NDCG with relevance labels from human judges and the Umbrela evaluator. We see that especially among top ranked systems, the evaluation strategy no longer agrees with human judges on which system is better. Axis ranges are adjusted to display the same top 60 (of 75) systems as in Figure 2.
4.1. Empirical Demonstration of the Risk
@@
The run circled in green in Figure 1 exemplifies this vulnerability. Submitted by team WaterlooClarke as run uwc1, this submission was deliberately designed to subvert the automatic evaluation process. Specifically, the team pooled the top 20 documents from 15 preliminary runs, spanning neural and traditional rankers, with and without query expansion. This pool was then judged by GPT-4o with the prompt described by Faggioli et al. (2023). Top-graded passages were subsequently judged pairwise following the procedure of Clarke et al. (2021) but substituting LLM-based assessments for crowdsourced human judgments.
@@
The final ranking for uwc1 was determined using LLM-based preference judgments as the primary key, LLM-based relevance assessments as the secondary key, and the reciprocal rank fusion (Cormack et al., 2009) of preliminary runs as a tertiary tie-breaker. As intended, this deliberate attempt to manipulate the evaluation process led uwc1 to rank significantly higher under automatic assessment (5th) than under manual assessment (28th).
4.2. LLM-based Relevance Assessment for Re-Ranking
@@
As Soboroff (2025) recognizes: “Retrieval and evaluation are the same problem. Asking a computer to decide if a document is relevant is no different than using a computer to retrieve documents and rank them in order of predicted degree of relevance.” In this sense, LLM-based relevance assessment can be viewed as a specific form of LLM-based re-ranking. To employ an LLM-based relevance assessment tool as a LLM-based re-ranker, one starts with an initial ranking and prompts the LLM to assign a score — expressed as a relevance grade — to each of the top-$k$ passages. These passages are then re-ranked according to their assigned relevance grades, preserving their original order among passages with equal grades.
@@
This re-ranking interpretation is further illustrated by another run submitted by team WaterlooClarke, uwc2. For uwc2, the team re-ranked the track’s baseline run using the prompt described by Arabzadeh and Clarke (2024). This re-ranking improves the baseline’s performance from 9th to 4th place under manual assessments, and from 7th to 3rd place under automatic assessments.
4.3. Simulation of Circularity
To explore the potential effects of widespread adoption, we simulate what would have happened if TREC RAG 2024 systems had incorporated Umbrela as part of their pipeline. Following the methodology described in Section 4.2, we apply Umbrela as a final-stage re-ranker to the outputs of all submitted retrieval systems. For this experiment, we re-rank using the Umbrela relevance labels from the track itself. While the track organizers employed Umbrela for evaluation, the labels it produced could instead have been employed as a final-stage ranker. While we use the labels from the track organizers, any track participant could just as easily have employed Umbrela as a final-stage re-ranker themselves.
Comparing original and re-ranked systems on manual (i.e., human) judgments in Figure 3, we see that re-ranking with the Umbrela judgments consistently improves system performance. TREC and similar evaluation experiments generally allow participants to use any automatic re-ranking process for their submissions. Since Umbrela re-ranking is an entirely automatic process, it immediately loses its value as a measurement tool for those ranking experiments. If any automatic re-ranking process can be used, using the measurement tool itself provides an optimal re-ranking (Soboroff, 2025). More generally, if system performance is measured solely by LLM-based tools, system developers are strongly incentivized to incorporate the same tools into their systems.
Once we adopt Umbrela as both a system component and an evaluation metric, it leads to an invalid circular evaluation. This effect is demonstrated in Figure 4, where the Umbrela-re-ranked runs from Figure 3 are evaluated with the same Umbrela relevance labels used for re-ranking and compared against evaluations based on human judgments. We observe a substantial increase in disagreement between the two evaluation methods, with discordant system pairs rising to 18% within the top 60 systems. As a result, Kendall’s $\tau$ drops sharply to 0.63. This degradation becomes even more pronounced at the top of the leaderboard: Kendall’s $\tau$ further decreases to 0.44 among the top 20 systems, 0.49 among the top 15, 0.38 among the top 10, and even turns negative ($\tau=-0.40$) among the top 5 systems. Under Umbrela-based evaluation, twelve systems now obtain NDCG scores exceeding 0.95, implying near-perfect ranking performance. Yet the same systems achieve manual NDCG scores only between 0.68 and 0.72, illustrating substantial score inflation due to circularity.
For publication in peer-reviewed information retrieval research venues it is often necessary to demonstrate that a proposed system significantly outperforms all strong baselines. Under a circular evaluation, however, such findings would no longer be credible. These results highlight the risks of using LLM-based evaluation pipelines without safeguards against feedback loops, especially when test collections are intended for reuse. Taking manually created relevance labels as the gold standard, we conclude that evaluating systems using the Umbrela LLM assessor — when those systems internally apply Umbrela-based re-ranking — results in an invalid circular experimental evaluation.
5. Automatic Judgments are not Gold Standards
@@
LLM-based relevance assessments cannot serve as a gold standard because they lack the grounding of a human carrying out an information task necessary to evaluate the usefulness of retrieval systems. A true gold standard must originate from human assessments, as only humans can determine the relevance of information in a way that reflects real-world utility.
@@
Faggioli et al. (2023) raised concerns about the potential unknown biases inherent in LLM-based assessments. However, one clear and concerning bias is that LLM-based relevance assessments tend to favor LLM-based ranking systems. Recently, Balog et al. (2025) report a detailed evaluation of how LLM-based rankers can influence LLM-based judges, providing the first empirical evidence that LLM judges exhibit “a clear and substantial bias in favor of LLM-based rankers.” This bias has been observed in other contexts as well (Liu et al., 2023; Wang et al., 2023; Panickssery et al., 2024), where LLMs demonstrate a form of “narcissism,” disproportionately favoring outputs generated by similar models. Furthermore, Alaofi et al. (2024) show that LLMs can be deceived through well-crafted prompt attacks embedded in content, leading them to incorrectly judge irrelevant text as relevant. These vulnerabilities highlight not only the susceptibility of LLM-based assessments to manipulation but also their inability to objectively evaluate diverse ranking approaches. Such biases and flaws further undermine the reliability of LLM-based assessments as a substitute for human judgments in critical tasks.
6. When Automatic Judgments become Useless